OpenStudy (anonymous):

Luke has a compost pit in his garden. The pit is 6 feet long, 3 feet wide, and 4 feet deep. She plans to make it 1.5 feet deeper. What will be the increase in the volume of the pit? Answer 27 cubic feet 36 cubic feet 99 cubic feet 108 cubic feet

OpenStudy (anonymous):

why rnt u getting it????/

OpenStudy (anonymous):

Volume = length * breadth * depth Earlier Volume Volume =9 * 6 * 2 = 108 feet^3 Now depth is increased by 1.5 feet so New Volume = 9 * 6 *3.5 =189feet^3 Increase Volume = 189 - 108 = 81 feet^3 Hope this helps! :D

OpenStudy (anonymous):

yes thank-you!! =)